OpenStudy (anonymous):

x - 2 and 5/7 = 2/3 Answer Options I got: (A) 2 and 1/21 (B) 2 and 8/21 (C) 3 and 1/3 (D) 3 and 8/21 I'm so stumped.

OpenStudy (anonymous):

i got A

OpenStudy (radar):

Why not D

OpenStudy (anonymous):

x-2+(5/7)=2/3 x-2(7/7)+(5/7)=2/3 x-(14/7)+(5/7)=2/3 x+(5-14)/7=2/3 x-(9/7)=2/3 x+(9/7)-(9/7)=(2/3)+(9/7) x=(2/3)(7/7)+(9/7)(3/3) x=(14/21)+(27/21) x=(14+27)/21 x=41/21 x=2-1/21

OpenStudy (anonymous):

2-(1/21)

OpenStudy (radar):

5 2 x - 2 -- = -- 7 3

OpenStudy (anonymous):

ahh i did it wrong so yea its whatever radar said

OpenStudy (anonymous):

sorry i thought the and meant '+' my bad

OpenStudy (anonymous):

Thanks for the help anyways:D

OpenStudy (radar):

I am trying to figure where jahtoday erred. His work looks good at least as far as I have followed it (down to the 3rd line)

OpenStudy (anonymous):

i did x-2+5/7 instead of x-2-5/7 also i ddnt get a valid answer choice

OpenStudy (radar):

Oh I see it at the beginning. Yes a sign problem! It was laid out nice. step by step.

OpenStudy (anonymous):

@ jahtoday \[\frac{41}{21}=1+\frac{20}{21} \]

OpenStudy (anonymous):

or 2-1/21

OpenStudy (anonymous):

2-1/21=42/21-1/21=41/21

OpenStudy (anonymous):

This is the form I intended.\[1\frac{20}{21} \]Pronounced, "One and twenty twenty- firsts." Never encountered the word "and" in the context of this problems's answer list to mean subtract.